OpenStudy (anonymous):

I Can not figure out how to do this one, any suggestions. \frac{ 1 }{ x }-\frac{ 1 }{ x-1}=\frac{ x-5 }{ 3x }

OpenStudy (anonymous):

\[ \frac{ 1 }{ x }-\frac{ 1 }{ x-1}=\frac{ x-5 }{ 3x }\]